KKK is currently unlevered and is valued at $5 million. The company is considering including debt in its capital structure and would like to know the likely impact on its value and cost of capital. The current cost of equity is 12 percent. The firm is considering offering $2 million of new debt with an interest rate of 7 percent. The KKK will use the proceeds of the debt issue to repurchase stock. There are currently 1,000,000 shares outstanding and the marginal tax rate is 34 percent. What are the new value and the new WACC for the KKK assuming that all cash flows are permanent?
Answer #1
Value of levered firm = Value of unlevered firm + debt*tax rate

=5+2*0.34= 5.68m

Equity = value -debt = 5.68-2= 3.68

D/E = 2/3.68=0.54347

Levered cost of equity = Unlevered cost of equity+D/E*( Unlevered cost of equity-cost of debt)*(1-tax rate)
Levered cost of equity = 12+0.54347*(12-7)*(1-0.34)
Levered cost of equity = 13.79
D/A = D/(E+D)
D/A = 0.54347/(1+0.54347)
=0.3521
Weight of equity = 1-D/A
Weight of equity = 1-0.3521
W(E)=0.6479
Weight of debt = D/A
Weight of debt = 0.3521
W(D)=0.3521
After tax cost of debt = cost of debt*(1-tax rate)
After tax cost of debt = 7*(1-0.34)
= 4.62
WACC=after tax cost of debt*W(D)+cost of equity*W(E)
WACC=4.62*0.3521+13.79*0.6479
WACC =10.56%